OpenStudy (anonymous):

The equation of line GH is y = 4x − 3. Write an equation of a line perpendicular to line GH in slope-intercept form that contains point (2, 3).

OpenStudy (anonymous):

y = 4x − 3 y = −4x − 5 11 over 4x − 3 y = negative 1 over 4x + 7 over 2

OpenStudy (anonymous):

. y = 4x − 3 y = −4x − 5 11 over 4x − 3 y = negative 1 over 4x + 7 over 2

OpenStudy (anonymous):

@Albany_Goon can you help please

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

hey :)

OpenStudy (anonymous):

hey lemme c here

OpenStudy (anonymous):

okay ^.^

OpenStudy (anonymous):

step 1: Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : y-(4*x-3)=0

OpenStudy (anonymous):

then Solve

OpenStudy (anonymous):

y-4x+3 = 0

OpenStudy (anonymous):

then

OpenStudy (anonymous):

im calculating lol

OpenStudy (anonymous):

ok but I recognize that we have here an equation of a straight line. Such an equation is usually written y=mx+b

OpenStudy (anonymous):

wat did you get nikki

OpenStudy (anonymous):

well I did the answers on a grid and on the last one I got a perpendicular line

OpenStudy (anonymous):



OpenStudy (anonymous):

lol somthimn like dat

OpenStudy (anonymous):

lol so I'm right?

OpenStudy (anonymous):

lol sure but Notice that when x = 0 the value of y is -3/1 so this line "cuts" the y axis at y=-3.0

OpenStudy (anonymous):

y-intercept = -3/1 = -3.0

OpenStudy (anonymous):

N When y = 0 the value of x is 3/4 Our line therefore "cuts" the x axis at x= 0.75000 x-intercept = 3/4 = 0.75

OpenStudy (anonymous):

Nikki?